A teacher has 10 boys and 12 girls in her grade school class. In how many ways can she select 9 of the children to be in a play if she must select 4 boys and 5 girls.
Answer by sudhanshu_kmr(1152)


number of ways = 10C4 *12C5 = 210*792 = 166320